# Low-temperature scaling laws in unconventional flat-band superconductors

### v1

@@ IV.3 Tunneling conductance
Under the assumption that the DOS of the superconductor in its normal state $D_{n}$ is energy independent, the tunneling current between a normal conductor and the superconductor is given by [16]
$I_{sn}=\frac{G_{nn}}{eD_{n}}\int_{-\infty}^{\infty}\mathrm{d}ED(E)\left[n_{\mathrm{F}}(E)-n_{\mathrm{F}}(E+eV)\right]\,$, (75)
where $G_{nn}$ is the differential conductance if the superconductor is driven into the normal state and $V$ is the bias voltage. The linear conductance is given by the derivative of the tunneling current with respect to the bias voltage at $V=0$,
$G_{sn}$ $=\frac{G_{nn}}{D_{n}}\int_{-\infty}^{\infty}\mathrm{d}x\,\frac{e^{x}}{(e^{x}+1)^{2}}D(xT)\,.$ (76)
@@
$G_{sn}(0)$ $=\frac{2G_{nn}}{D_{n}}\Gamma(\alpha)\eta(\alpha-1)T^{\alpha-1}\ln^{\beta}(1/T)\,$, (77)
@@ V Application to $C_{6v}$-symmetric systems
The hexagonal group $C_{6v}$ possesses six irreducible representations, where three of them are even-parity representations (spin-singlet states) and the other three are odd-parity representations (spin-triplet states) [52, 53]. Table 3 presents the basis function of each irreducible representation together with the corresponding low-temperature scaling laws. For all superconducting states considered in Table 3, we can find parameters $m$, $L$, and $q$ such that the basis functions can be cast into the dispersion given in Eq. (18).
@@ VI Summary and discussion
We calculated the low-temperature scaling laws of the order parameter, the superfluid weight, the tunneling conductance, the specific heat, the Sommerfeld coefficient, and the spin-lattice relaxation rate in flat-band superconductors with unconventional pairing. The results obtained for different node types are collected in Table 1. In particular, by comparing the scaling laws of the geometrical and functional contributions to the superfluid weight, we found the functional superfluid weight to be subleading for the nodal structures considered here. Therefore, we conclude that the scaling laws of the superfluid weight reported in Ref. [24] are intact.
Several generalizations of our work are possible. First of all, we mostly considered the dispersion given in Eq. (18) which is a special case of the dispersion in Eq. (15). It would be then interesting to further explore nodal structures not covered here but still allowed by the Weierstrass preparation theorem. For example, node types that exhibit “cusps” such as the dispersion in Eq. (16) may appear in non-centrosymmetric crystals [57], and probably give rise to different exponents or additional logarithms in the scaling laws via the Puiseux expansion. Also, it remains to be clarified to what extent these scaling laws carry over to quasicrystals [58], and how the temperature dependence changes near the critical temperature [25].
@@ Acknowledgements.
Appendix A Functional superfluid weight for nonzero temperatures
In this section, we derive an expression for the functional superfluid weight for nonzero temperature [23],
$D_{\mathrm{s},ij}^{\mathrm{func}}(T)$ $\coloneqq\frac{1}{V}\sum_{\mathbf{k},\mathbf{k}^{\prime}}S_{i,\alpha\mu}(\mathbf{k})M^{-1}_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k}^{\prime})S_{j,\beta\nu}(\mathbf{k}^{\prime})\,$, (84)
@@ Appendix A Functional superfluid weight for nonzero temperatures
$S_{i\alpha\mu}(\mathbf{k})$ $\coloneqq\frac{\delta}{\delta\Delta^{\mu}_{\alpha}(\mathbf{q};\mathbf{k})}\left(\frac{\partial\Omega}{\partial q_{i}}\right)\biggl|_{\mathbf{q}=0}\,,\qquad M_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k}^{\prime})\coloneqq\frac{\delta^{2}\Omega}{\delta\Delta^{\mu}_{\alpha}(\mathbf{q};\mathbf{k})\delta\Delta^{\nu}_{\beta}(\mathbf{q};\mathbf{k}^{\prime})}\biggl|_{\mathbf{q}=0}\,$, (85)
and $\Omega$ is the grand potential of mean-field BCS theory given in Eq. (1). According to the chain rule, the Hessian matrix $M$ is obtained as
$M_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k}^{\prime})=\sum_{\mathbf{k}^{\prime\prime},n}\Bigg[n_{\mathrm{F}}^{\prime}(E_{\mathbf{k}^{\prime\prime}n}(\mathbf{q}))\frac{\delta E_{\mathbf{k}^{\prime\prime}n}(\mathbf{q})}{\delta\Delta^{\mu}_{\alpha}(\mathbf{q};\mathbf{k})}\frac{\delta E_{\mathbf{k}^{\prime\prime}n}(\mathbf{q})}{\delta\Delta^{\nu}_{\beta}(\mathbf{q};\mathbf{k}^{\prime})}+n_{\mathrm{F}}(E_{\mathbf{k}^{\prime\prime}n}(\mathbf{q}))\frac{\delta^{2}E_{\mathbf{k}^{\prime\prime}n}(\mathbf{q})}{\delta\Delta^{\mu}_{\alpha}(\mathbf{q};\mathbf{k})\delta\Delta^{\nu}_{\beta}(\mathbf{q};\mathbf{k}^{\prime})}\Bigg]+VU^{-1}(\mathbf{k},\mathbf{k}^{\prime})\delta_{\alpha\beta}\delta_{\mu\nu}\,.$ (86)
To calculate the derivatives of the energies, we make use of the Hellman-Feynman theorem [59, 60], which states
@@
$\frac{\delta\mathcal{H}_{\mathrm{BdG}}(\mathbf{k}^{\prime\prime},\mathbf{q})}{\delta\Delta^{\mu}_{\alpha}(\mathbf{q};\mathbf{k})}\biggl|_{\mathbf{q}=0}$ $=\begin{pmatrix}A^{(0)}_{\alpha\mu}(\mathbf{k})&A_{\alpha\mu}(\mathbf{k})\\ A^{\dagger}_{\alpha\mu}(\mathbf{k})&A^{(1)}_{\alpha\mu}(\mathbf{k})\end{pmatrix}\delta_{\mathbf{k},\mathbf{k}^{\prime\prime}}$ (89)
with
$A^{(0)}_{\alpha\mu}(\mathbf{k})$ $=\mathcal{V}_{-\mathbf{k}}^{T}\Lambda_{\alpha\mu}(\mathbf{k})\mathcal{U}_{\mathbf{k}}+\mathcal{U}_{\mathbf{k}}^{\dagger}\Lambda_{\alpha\mu}^{\dagger}(\mathbf{k})\mathcal{V}^{\ast}_{-\mathbf{k}}\,$, (90)
$A^{(1)}_{\alpha\mu}(\mathbf{k})$ $=\mathcal{U}_{-\mathbf{k}}^{T}\Lambda_{\alpha\mu}(\mathbf{k})\mathcal{V}_{\mathbf{k}}+\mathcal{V}_{\mathbf{k}}^{\dagger}\Lambda_{\alpha\mu}^{\dagger}(\mathbf{k})\mathcal{U}^{\ast}_{-\mathbf{k}}\,$, (91)
$A_{\alpha\mu}(\mathbf{k})$ $=\mathcal{V}_{-\mathbf{k}}^{T}\Lambda_{\alpha\mu}(\mathbf{k})\mathcal{V}_{\mathbf{k}}+\mathcal{U}_{\mathbf{k}}^{\dagger}\Lambda_{\alpha\mu}^{\dagger}(\mathbf{k})\mathcal{U}^{\ast}_{-\mathbf{k}}\,$, (92)
$\Lambda_{\alpha\mu}(\mathbf{k})$ $=\begin{cases}\phantom{i}S^{\dagger}(\mathbf{k})\mathbbm{1}_{\alpha}S(\mathbf{k})&\mu=\mathrm{R}\,,\\ iS^{\dagger}(\mathbf{k})\mathbbm{1}_{\alpha}S(\mathbf{k})&\mu=\mathrm{I}\,.\end{cases}$ (93)
Therefore, the Hessian matrix is given by
$M_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k}^{\prime})=$ $VU^{-1}(\mathbf{k},\mathbf{k}^{\prime})\delta_{\alpha\beta}\delta_{\mu\nu}-\Pi_{\alpha\mu,\beta\nu}(\mathbf{k})\delta_{\mathbf{k},\mathbf{k}^{\prime}}$ (94)
with
$\Pi_{\alpha\mu,\beta\nu}(\mathbf{k})$ $=\sum_{\gamma,\gamma^{\prime}=1}^{N_{\mathrm{B}}}\Bigg[\frac{1-n_{\mathrm{F}}(E_{\mathbf{k}\gamma})-n_{\mathrm{F}}(E_{-\mathbf{k}\gamma^{\prime}})}{E_{\mathbf{k}\gamma}+E_{-\mathbf{k}\gamma^{\prime}}}\mathrm{Re}([A_{\alpha\mu}(\mathbf{k})]_{\gamma\gamma^{\prime}}[A^{\dagger}_{\beta\nu}(\mathbf{k})]_{\gamma^{\prime}\gamma})+(\alpha\mu\gamma\leftrightarrow\beta\nu\gamma^{\prime})$
$-\frac{n_{\mathrm{F}}(E_{\mathbf{k}\gamma})-n_{\mathrm{F}}(E_{\mathbf{k}\gamma^{\prime}})}{E_{\mathbf{k}\gamma}-E_{\mathbf{k}\gamma^{\prime}}}\mathrm{Re}([A^{(0)}_{\alpha\mu}(\mathbf{k})]_{\gamma\gamma^{\prime}}[A^{(0)}_{\beta\nu}(\mathbf{k})]_{\gamma^{\prime}\gamma})-\frac{n_{\mathrm{F}}(E_{-\mathbf{k}\gamma^{\prime}})-n_{\mathrm{F}}(E_{-\mathbf{k}\gamma})}{E_{-\mathbf{k}\gamma^{\prime}}-E_{-\mathbf{k}\gamma}}\mathrm{Re}([A^{(1)}_{\alpha\mu}(\mathbf{k})]_{\gamma\gamma^{\prime}}[A^{(1)}_{\beta\nu}(\mathbf{k})]_{\gamma^{\prime}\gamma})\Bigg]\,.$ (95)
Here, we used the property $E_{\mathbf{k}n}=-E_{-\mathbf{k}(n+N_{\mathrm{B}})}$ of the eigenvalues and the fact that the Fermi distribution function satisfies $n_{\mathrm{F}}(-E_{\mathbf{k}n})=1-n_{\mathrm{F}}(E_{\mathbf{k}n})$. Moreover, the prefactors in the summations of Eq. (95) should be understood as $n_{\mathrm{F}}^{\prime}(E_{\mathbf{k}\gamma})$ when $E_{\mathbf{k}\gamma}=E_{\mathbf{k}\gamma^{\prime}}$ [21]. The inverse of $M$ corresponds to a geometric series,
$M^{-1}_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k^{\prime}})$ $=\frac{1}{V}U(\mathbf{k},\mathbf{k}^{\prime})\delta_{\alpha\beta}\delta_{\mu\nu}+\frac{1}{V}\sum_{n=1}^{\infty}\frac{1}{V^{n}}\sum_{\begin{subarray}{c}\mu_{1},\ldots,\mu_{n}=\mathrm{R},\mathrm{I}\\ \mathbf{k}_{1},\ldots,\mathbf{k}_{n}\\ \alpha_{1},\ldots,\alpha_{n}\end{subarray}}U(\mathbf{k},\mathbf{k}_{1})\Pi_{\alpha_{1}\mu_{1},\alpha_{2}\mu_{2}}(\mathbf{k}_{1})U(\mathbf{k}_{1},\mathbf{k}_{2})\ldots\Pi_{\alpha_{n-1}\mu_{n-1},\alpha_{n}\mu_{n}}(\mathbf{k}_{n})U(\mathbf{k}_{n},\mathbf{k}^{\prime})$ (96)
$=\frac{U_{0}}{V}\sum_{i,j=1}^{d_{\Gamma}}\psi_{i}^{\Gamma}(\mathbf{k})\bar{\psi}_{j}^{\Gamma}(\mathbf{k}^{\prime})\left[(\mathbbm{1}+\Sigma)^{-1}\right]_{i\alpha\mu,j\beta\nu}\,$, (97)
where we applied the Sherman-Morrison-Woodbury formula in the second line for the pairing potential (7) (see also Ref. [23] for more details) and the ($2d_{\Gamma}N_{\mathrm{B}}\times 2d_{\Gamma}N_{\mathrm{B}}$)-matrix $\Sigma$ is defined via
$\Sigma_{i\alpha\mu,j\beta\nu}=-\frac{U_{0}}{V}\sum_{\mathbf{k}}\psi_{i}^{\Gamma}(\mathbf{k})\bar{\psi}_{j}^{\Gamma}(\mathbf{k})\Pi_{\alpha\mu,\beta\nu}(\mathbf{k})\,.$ (98)
We proceed analogously to calculate the functional derivative of $\partial\Omega/\partial q_{i}$. According to the Hellmann-Feynman theorem, we obtain
@@
Note that the functional derivative of the partial derivative of the BdG Hamiltonian is zero. The first factor in Eq. (100) was already calculated in Eq. (89). The second factor can be calculated similarly as
[…]
with
$B_{i}^{(0)}(\mathbf{k})$ $=\mathcal{U}^{\dagger}_{\mathbf{k}}\Xi_{i}(\mathbf{k})\mathcal{U}_{\mathbf{k}}+\mathcal{V}^{T}_{-\mathbf{k}}\Xi_{i}(\mathbf{k})\mathcal{V}^{\ast}_{-\mathbf{k}}\,$, (102)
$B_{i}^{(1)}(\mathbf{k})$ $=\mathcal{V}^{\dagger}_{\mathbf{k}}\Xi_{i}(\mathbf{k})\mathcal{V}_{\mathbf{k}}+\mathcal{U}^{T}_{-\mathbf{k}}\Xi_{i}(\mathbf{k})\mathcal{U}^{\ast}_{-\mathbf{k}}\,$, (103)
$B_{i}(\mathbf{k})$ $=\mathcal{U}^{\dagger}_{\mathbf{k}}\Xi_{i}(\mathbf{k})\mathcal{V}_{\mathbf{k}}+\mathcal{V}^{T}_{-\mathbf{k}}\Xi_{i}(\mathbf{k})\mathcal{U}^{\ast}_{-\mathbf{k}}\,$, (104)
$\Xi_{i}(\mathbf{k})$ $=S^{\dagger}(\mathbf{k})\partial_{k_{i}}H(\mathbf{k})S(\mathbf{k})\,.$ (105)
Therefore, we obtain
$S_{i,\alpha\mu}(\mathbf{k})$ $=\sum_{\gamma,\gamma^{\prime}=1}^{N_{\mathrm{B}}}\Bigg[\frac{1-n_{\mathrm{F}}(E_{\mathbf{k}\gamma})-n_{\mathrm{F}}(E_{-\mathbf{k}\gamma^{\prime}})}{E_{\mathbf{k}\gamma}+E_{-\mathbf{k}\gamma^{\prime}}}\mathrm{Re}([A_{\alpha\mu}(\mathbf{k})]_{\gamma\gamma^{\prime}}[B^{\dagger}_{i}(\mathbf{k})]_{\gamma^{\prime}\gamma})+\frac{1-n_{\mathrm{F}}(E_{-\mathbf{k}\gamma})-n_{\mathrm{F}}(E_{\mathbf{k}\gamma^{\prime}})}{E_{-\mathbf{k}\gamma}+E_{\mathbf{k}\gamma^{\prime}}}\mathrm{Re}([A^{\dagger}_{\alpha\mu}(\mathbf{k})]_{\gamma\gamma^{\prime}}[B_{i}(\mathbf{k})]_{\gamma^{\prime}\gamma})$
$-\frac{n_{\mathrm{F}}(E_{\mathbf{k}\gamma})-n_{\mathrm{F}}(E_{\mathbf{k}\gamma^{\prime}})}{E_{\mathbf{k}\gamma}-E_{\mathbf{k}\gamma^{\prime}}}\mathrm{Re}([A^{(0)}_{\alpha\mu}(\mathbf{k})]_{\gamma\gamma^{\prime}}[B^{(0)}_{i}(\mathbf{k})]_{\gamma^{\prime}\gamma})-\frac{n_{\mathrm{F}}(E_{-\mathbf{k}\gamma^{\prime}})-n_{\mathrm{F}}(E_{-\mathbf{k}\gamma})}{E_{-\mathbf{k}\gamma^{\prime}}-E_{-\mathbf{k}\gamma}}\mathrm{Re}([A^{(1)}_{\alpha\mu}(\mathbf{k})]_{\gamma\gamma^{\prime}}[B^{(1)}_{i}(\mathbf{k})]_{\gamma^{\prime}\gamma})\Bigg]\,.$ (106)
Appendix B Superfluid weight of isolated flat-band superconductors
For the discussion of the low-temperature behavior of the functional superfluid weight, the general expression is not necessary and we can work in the isolated-band limit. Throughout this calculation, we assume that the gap function is proportional to the identity matrix (uniform pairing condition) and that the band with index $n_{0}$ is flat and located at the chemical potential.
In the case of isolated bands at the zeroth-order perturbation theory, it has been shown that the Bogoliubov eigenvalues of the $n$th band are given by
$E_{\mathbf{k}n}^{(0)}(\mathbf{q})=\sqrt{(\varepsilon_{n}(\mathbf{k}\pm\mathbf{q})-\mu)^{2}+\mathcal{D}_{n,\mathbf{k}}(\mathbf{q})\mathcal{D}_{n,\mathbf{k}}^{\dagger}(\mathbf{q})}$ (107)
[…]
$\frac{\partial^{2}}{\partial q_{i}\partial q_{j}}\left(-T\ln(1+\exp\left(-\frac{E_{\mathbf{k}n}^{(0)}(\mathbf{q})}{T}\right))\right)\biggl|_{\mathbf{q}=0}$ $=n_{\mathrm{F}}^{\prime}(E_{\mathbf{k}n}^{(0)}(\mathbf{q}))\frac{\partial E_{\mathbf{k}n}^{(0)}(\mathbf{q})}{\partial q_{i}}\frac{\partial E_{\mathbf{k}n}^{(0)}(\mathbf{q})}{\partial q_{j}}+n_{\mathrm{F}}(E_{\mathbf{k}n}^{(0)}(\mathbf{q}))\frac{\partial^{2}E_{\mathbf{k}n}^{(0)}(\mathbf{q})}{\partial q_{i}\partial q_{j}}\biggl|_{\mathbf{q}=0}$ (108)
$\approx\frac{n_{\mathrm{F}}(E_{\mathbf{k}n}^{(0)})}{E_{\mathbf{k}n}^{(0)}}|f(\mathbf{k})|^{2}g_{ij}^{(n)}(\mathbf{k})\,$, (109)
where we used $\partial_{q_{i}}(\mathcal{D}_{n,\mathbf{k}}(\mathbf{q})\mathcal{D}_{n,\mathbf{k}}^{\dagger}(\mathbf{q}))_{\mathbf{q}=0}=0$ due to the uniform pairing condition. We can neglect the curvature of the other nonflat bands because they are well separated from the flat band such that $E_{\mathbf{k}n}^{(0)}\gg E_{\mathbf{k}}\equiv E_{\mathbf{k}n_{0}}^{(0)}$ for $n\neq n_{0}$. Therefore, we can approximate the geometrical superfluid weight by
$D_{\mathrm{s},ij}^{\mathrm{geom}}(T)\approx\frac{1}{V}\sum_{\mathbf{k}}\frac{n_{\mathrm{F}}(E_{\mathbf{k}})}{E_{\mathbf{k}}}|f(\mathbf{k})|^{2}g_{ij}(\mathbf{k})\,$, (110)
where $g_{ij}=g_{ij}^{(n_{0})}$ represents the quantum metric of the $n_{0}$th band. This expression is used to calculate the temperature dependence of the geometrical superfluid weight.
Analogously, using
$\frac{\delta}{\delta\Delta_{\alpha}^{\mu}(\mathbf{q};\mathbf{k}^{\prime})}\left(\frac{\partial}{\partial q_{i}}\left(-T\ln(1+\exp\left(-\frac{E_{\mathbf{k}n}^{(0)}(\mathbf{q})}{T}\right))\right)\right)\biggl|_{\mathbf{q}=0}$ $\approx-\frac{n_{\mathrm{F}}(E_{\mathbf{k}n}^{(0)}){f}^{\bar{\mu}}(\mathbf{k})}{2E_{\mathbf{k}n}^{(0)}}\mathrm{Im}(R^{(1)}_{n,\alpha,i}(\mathbf{k})+R^{(2)}_{n,\alpha,i}(\mathbf{k}))$ (111)
and following the calculation in Ref. [23], we obtain the following expression for the functional contribution,
$D_{\mathrm{s},ij}^{\mathrm{func}}\approx\frac{1}{V}\sum_{n,n^{\prime},\mathbf{k},\mathbf{k}^{\prime},\alpha,\beta,\mu,\nu}\frac{M_{\alpha\mu,\beta\nu}^{-1}(\mathbf{k},\mathbf{k}^{\prime})f^{\mu}(\mathbf{k})f^{\nu}(\mathbf{k}^{\prime})n_{\mathrm{F}}(E_{\mathbf{k}n}^{(0)})n_{\mathrm{F}}(E_{\mathbf{k}^{\prime}n^{\prime}}^{(0)})}{E_{\mathbf{k}n}^{(0)}E_{\mathbf{k}^{\prime}n^{\prime}}^{(0)}}h_{ij,\alpha\beta}^{(nn^{\prime})}(\mathbf{k},\mathbf{k}^{\prime})\,$, (112)
where

### v2

@@ IV.3 Tunneling conductance
The tunneling current between a superconductor and a normal conductor can be phenomenologically described by [51, §3.8]
$I_{sn}=A\int_{-\infty}^{\infty}\mathrm{d}ED(E)\left[n_{\mathrm{F}}(E)-n_{\mathrm{F}}(E+V)\right]\,$, (76)
where $V$ is the bias voltage and we work with the convention $e=1$ as indicated below Eq. (1). The proportionality coefficient $A$ depends on the tunneling matrix elements and the DOS of the normal conductor, both of which are assumed to be energy independent. This expression corresponds to the tunneling (low transparency) limit in which transport is characterized by the quasiparticle DOS. If the interface transparency is not small, or interface-induced bound states are relevant, Andreev reflection becomes important and the DOS-based description may no longer be sufficient. In that case, a treatment within the Blonder-Tinkham-Klapwijk (BTK) framework is more appropriate [7, 61].
The linear conductance is given by the derivative of the tunneling current with respect to the bias voltage at $V=0$,
$G_{sn}$ $=A\int_{-\infty}^{\infty}\mathrm{d}x\,\frac{e^{x}}{(e^{x}+1)^{2}}D(xT)\,.$ (77)
@@
$G_{sn}$ $=2AD_{0}\Gamma(\alpha)\eta(\alpha-1)T^{\alpha-1}\ln^{\beta}(1/T)\,$, (78)
@@ V Application to $C_{6v}$-symmetric systems
The hexagonal group $C_{6v}$ possesses six irreducible representations, where three of them are even-parity representations (spin-singlet states) and the other three are odd-parity representations (spin-triplet states) [42, 6]. Table 3 presents the basis function of each irreducible representation together with the corresponding low-temperature scaling laws. For all superconducting states considered in Table 3, we have $b=1$, and the other parameters $m$, $L$, and $q$ can be found such that the basis functions can be cast into the dispersion given in Eq. (18).
@@ VI Summary and discussion
We calculated the low-temperature scaling laws of the order parameter, the superfluid weight, the tunneling conductance, the specific heat, the Sommerfeld coefficient, and the spin-lattice relaxation rate in flat-band superconductors with unconventional pairing. The results obtained for different node types are collected in Table 1. In particular, by comparing the scaling laws of the geometrical and functional contributions to the superfluid weight, we found them to be of the same order for all one-dimensional superconducting states and most of the higher-dimensional superconducting states and conclude that the scaling laws of the superfluid weight reported in Ref. [20] are intact. However, once we have a superconducting state with $b<1$, the functional contribution might become dominant such that the scaling laws in Ref. [20] need to be adjusted.
Several generalizations of our work are possible. First of all, we mostly considered the dispersion given in Eq. (18) which is a special case of the dispersion in Eq. (15). It would be then interesting to further explore nodal structures not covered here but still allowed by the Weierstrass preparation theorem. For example, node types that exhibit cusps such as the dispersion in Eq. (16) may appear in non-centrosymmetric crystals [26], and probably give rise to different exponents or additional logarithms in the scaling laws via the Puiseux expansion. Also, it remains to be clarified to what extent these scaling laws carry over to quasicrystals [47], and how the temperature dependence changes near the critical temperature [40].
@@ Acknowledgements.
Data availability
No data were created or analyzed in this study.
Appendix A Superfluid weight of isolated flat-band superconductors for arbitrary temperature
In this appendix, we derive expressions for the superfluid weight in the isolated flat-band limit necessary for the discussion of the low-temperature behavior. While Ref. [9] derived the expressions only in the zero-temperature limit and assumed the presence of a real-valued pairing potential, here we derive the expressions for the superfluid weight of an isolated flat-band superconductor with a complex-valued pairing potential at arbitrary temperature. We follow the derivation presented in Ref. [9]. Throughout this calculation, we assume that the gap function is proportional to the identity matrix (uniform pairing condition) and that the band with index $n_{0}$ is flat and located at the chemical potential. As stated in Eq. (42), the superfluid weight consists of three contributions. These can be obtained via derivatives of the mean-field BCS grand potential $\Omega$, defined in Eq. (1), via
$D_{\mathrm{s},ij}^{\mathrm{conv}}(T)+D_{\mathrm{s},ij}^{\mathrm{geom}}(T)=\frac{1}{V}\frac{\partial^{2}\Omega}{\partial q_{i}\partial q_{j}}\biggl|_{\mathbf{q}=0}\,,\qquad D_{\mathrm{s},ij}^{\mathrm{func}}(T)$ $=\frac{1}{V}\sum_{\mathbf{k},\mathbf{k}^{\prime}}S_{i,\alpha\mu}(\mathbf{k})M^{-1}_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k}^{\prime})S_{j,\beta\nu}(\mathbf{k}^{\prime})\,$, (85)
@@ Appendix A Superfluid weight of isolated flat-band superconductors for arbitrary temperature
$S_{i\alpha\mu}(\mathbf{k})$ $\coloneqq\frac{\delta}{\delta f^{\mu}_{\alpha}(\mathbf{q};\mathbf{k})}\left(\frac{\partial\Omega}{\partial q_{i}}\right)\biggl|_{\mathbf{q}=0}\,,\qquad M_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k}^{\prime})\coloneqq\frac{\delta^{2}\Omega}{\delta f^{\mu}_{\alpha}(\mathbf{q};\mathbf{k})\delta f^{\nu}_{\beta}(\mathbf{q};\mathbf{k}^{\prime})}\biggl|_{\mathbf{q}=0}\,$, (86)
and $f^{\mathrm{R}}=\mathrm{Re}(f)$ and $f^{\mathrm{I}}=\mathrm{Im}(f)$ indicate the real and imaginary parts. In the case of isolated bands, due to TRS, the Bogoliubov eigenvalues of the $n$th band are given by
$E_{\mathbf{k}n\pm}^{(0)}(\mathbf{q})=\pm\sqrt{(\varepsilon_{n}(\mathbf{k})-\mu)^{2}+|\mathcal{D}_{n,\mathbf{k}}(\mathbf{q})|^{2}}+\ldots\,$, (87)
[…]
$\frac{\partial^{2}}{\partial q_{i}\partial q_{j}}\left(-T\ln(1+\exp\left(-\frac{E_{\mathbf{k}n\pm}^{(0)}(\mathbf{q})}{T}\right))\right)\Bigg|_{\mathbf{q}=0}$ $=n_{\mathrm{F}}^{\prime}(E_{\mathbf{k}n\pm}^{(0)}(\mathbf{q}))\frac{\partial E_{\mathbf{k}n\pm}^{(0)}(\mathbf{q})}{\partial q_{i}}\frac{\partial E_{\mathbf{k}n\pm}^{(0)}(\mathbf{q})}{\partial q_{j}}+n_{\mathrm{F}}(E_{\mathbf{k}n\pm}^{(0)}(\mathbf{q}))\frac{\partial^{2}E_{\mathbf{k}n\pm}^{(0)}(\mathbf{q})}{\partial q_{i}\partial q_{j}}\biggl|_{\mathbf{q}=0}$ (88)
$\approx-\frac{4n_{\mathrm{F}}(E_{\mathbf{k}n\pm}^{(0)})}{E_{\mathbf{k}n\pm}^{(0)}}|f(\mathbf{k})|^{2}g_{ij}^{(n)}(\mathbf{k})\,$, (89)
where we used $\partial_{q_{i}}|\mathcal{D}_{n,\mathbf{k}}(\mathbf{q})|^{2}|_{\mathbf{q}=0}=0$ due to the uniform pairing condition. We can neglect the curvature of the other nonflat bands because they are well separated from the flat band such that $E_{\mathbf{k}n}^{(0)}\gg E_{\mathbf{k}}\equiv E_{\mathbf{k}n_{0}}^{(0)}$ for $n\neq n_{0}$. By using $n_{\mathrm{F}}(-E)=1-n_{\mathrm{F}}(E)$, we can approximate the geometrical superfluid weight by
$D_{\mathrm{s},ij}^{\mathrm{geom}}(T)\approx\frac{4}{V}\sum_{\mathbf{k}}\frac{{1-2n_{\mathrm{F}}(E_{\mathbf{k}})}}{E_{\mathbf{k}}}|f(\mathbf{k})|^{2}g_{ij}(\mathbf{k})\,$, (90)
where $g_{ij}\equiv g_{ij}^{(n_{0})}$ represents the quantum metric of the $n_{0}$th band [43]. This expression is used to calculate the temperature dependence of the geometrical superfluid weight. Analogously, we have
@@
$D_{\mathrm{s},ij}^{\mathrm{func}}(T)\approx\frac{4}{V}\sum_{\mathbf{k},\mathbf{k}^{\prime},\alpha,\beta,\mu,\nu}\frac{M_{\alpha\mu,\beta\nu}^{-1}(\mathbf{k},\mathbf{k}^{\prime}){s_{\mu}s_{\nu}}f^{\bar{\mu}}(\mathbf{k})f^{\bar{\nu}}(\mathbf{k}^{\prime}){(2n_{\mathrm{F}}(E_{\mathbf{k}})-1)(2n_{\mathrm{F}}(E_{\mathbf{k}^{\prime}})-1)}}{E_{\mathbf{k}}E_{\mathbf{k}^{\prime}}}h_{ij,\alpha\beta}(\mathbf{k},\mathbf{k}^{\prime})\,$, (92)
where
$h_{ij,\alpha\beta}(\mathbf{k},\mathbf{k}^{\prime})=\frac{1}{2}\sum_{m,m^{\prime}\neq n_{0}}\mathrm{Re}\!\left[O^{(1)}_{n_{0}n_{0}mm^{\prime},\alpha\beta}(\mathbf{k},\mathbf{k}^{\prime})e^{(n_{0})}_{i,m}(\mathbf{k})\bar{e}^{(n_{0})}_{j,m^{\prime}}(\mathbf{k}^{\prime})+O^{(2)}_{n_{0}n_{0}mm^{\prime},\alpha\beta}(\mathbf{k},\mathbf{k}^{\prime})e^{(n)}_{i,m}(\mathbf{k})\bar{e}^{(m^{\prime})}_{j,n_{0}}(\mathbf{k}^{\prime})\right]$ (93)
[…]
$M_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k}^{\prime})=\sum_{\mathbf{k}^{\prime\prime},n,\sigma}\Bigg[n_{\mathrm{F}}^{\prime}(E_{\mathbf{k}^{\prime\prime}n\sigma}^{(0)}(\mathbf{q}))\frac{\delta E_{\mathbf{k}^{\prime\prime}n\sigma}^{(0)}(\mathbf{q})}{\delta f^{\mu}_{\alpha}(\mathbf{q};\mathbf{k})}\frac{\delta E_{\mathbf{k}^{\prime\prime}n\sigma}^{(0)}(\mathbf{q})}{\delta f^{\nu}_{\beta}(\mathbf{q};\mathbf{k}^{\prime})}+n_{\mathrm{F}}(E_{\mathbf{k}^{\prime\prime}n\sigma}^{(0)}(\mathbf{q}))\frac{\delta^{2}E_{\mathbf{k}^{\prime\prime}n\sigma}^{(0)}(\mathbf{q})}{\delta f^{\mu}_{\alpha}(\mathbf{q};\mathbf{k})\delta f^{\nu}_{\beta}(\mathbf{q};\mathbf{k}^{\prime})}\Bigg]\Bigg|_{\mathbf{q}=0}+VU_{\mu\nu}^{-1}(\mathbf{k},\mathbf{k}^{\prime})\delta_{\alpha\beta}\,$, (94)
where $U_{\mu\nu}(\mathbf{k},\mathbf{k}^{\prime})=\mathrm{Re}(U(\mathbf{k},\mathbf{k}^{\prime}))\delta_{\mu\nu}-\mathrm{Im}(U(\mathbf{k},\mathbf{k}^{\prime}))\epsilon_{\mu\nu}$. In particular, by making use of
@@
and defining
@@
with the Sherman-Morrison-Woodbury formula [56], we find the inverse of $M$ to correspond to a geometric series as
$M^{-1}_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k^{\prime}})=$ $\frac{1}{V}U_{\mu\nu}(\mathbf{k},\mathbf{k}^{\prime})\delta_{\alpha\beta}$